OpenStudy (anonymous):

Evaluate 8 m plus nineteen minus m when m equals 3. forty ninety-nine one hundred sixty-five one hundred seventy-three

OpenStudy (anonymous):

hey bro

OpenStudy (anonymous):

40

OpenStudy (anonymous):

def 40

OpenStudy (gorv):

8m+19-m at m=3 24+19-3=40

OpenStudy (anonymous):

8(3)+19-3=40

OpenStudy (anonymous):

thx to everyone